# Optimistic Online Non-stochastic Control via FTRL

### v1

@@ V Numerical Example
Recall that OptFTRL-C was designed to take advantage of a prediction oracle that forecasts future cost functions with unknown accuracy. Theorem 1 then demonstrated that the average policy regret $\nicefrac{{\mathcal{R}_{T}}}{{T}}$ of OptFTRL-C always converges to 00, but does so faster for accurate predictions. We therefore plot the policy regret of OptFTRL-C when provided with either accurate or inaccurate predictions. The implementation code of the policies OptFTRL-C, GPC, and the benchmark $\pi^{\star}$, as well as the code necessary to reproduce all experiments, is available at the GitHub repository [47].
@@
In the accurate prediction case, we set $\tilde{c}_{t}(\cdot,\cdot)=c_{t}(\cdot,\cdot)$ with probability $\rho=0.9$. Otherwise, we set $\tilde{c}_{t}(\cdot,\cdot)$ to be uniformally random ($\boldsymbol{\alpha}_{t}\in[-1,1]$). Hence, $\rho$ represents the probability of correctly predicting $c_{t}(\cdot,\cdot)$, and we sample it at every slot. For inaccurate prediction, we set $\rho=0.1$. We compare with GPC, which was shown to outperform the classical $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ controllers in a wide range of situations [^8] For additional details on the performance of GPC vs traditional ones, see the experiments in the tutorial [48], and its associated codebase [49]..
@@
In scenario (b), we deploy an alternating cost function. Namely, $\boldsymbol{\alpha}_{t}$ alternate between $(1,1)$ and $(-0.5,-0.5)$ every $50$ steps. The disturbances are still $\boldsymbol{w}_{t}\!=\!(1,1),\forall t$. This alternating cost represents an adversarial fluctuation in the cost trajectory, and it is where the non-stochastic framework demonstrates its efficacy. Namely, this fluctuation is enough to violate the guarantees of $\mathcal{H}_{2}$ controllers, but at the same time, it is small in magnitude, rendering $\mathcal{H}_{\infty}$ controller overly pessimistic. It is worth noting that the fluctuation observed in Fig. $b$ is attributed to the update rule of GPC. This update directly modifies the decision variables $M_{t+1}$ based on the observed cost. In contrast, FTRL aggregates both past and future costs to determine $M_{t+1}$. With accurate predictions, this method does not induce much fluctuation as it foresees the upcoming small disturbance. Overall, OptFTRL-C with good prediction achieves an improvement of $32\%$ in $\nicefrac{{\mathcal{R}_{T}}}{{T}}$ value over GPC, while having a $13.4\%$ degradation when fed with the inaccurate oracle.
In scenario (c) we also deploy an alternating cost function but with different lower magnitudes. Namely, $\boldsymbol{\alpha}_{t}$ alternate between $(0.1,0.1)$ and $(-0.5,-0.5)$ every $50$ steps, and $w_{t}=(0.1,0.1)$. The goal of this scenario is to show that OptFTRL-C can have an advantage over GPC even regardless of the prediction quality through adapatablity to “easy” environments (i.e., environments with small gradients). In general, GPC performance takes a hit since its learning rate is tuned with the upper bound for the gradient (i.e., $\alpha=w=\sqrt{2},T=10^{3}$). The alternating frequency, on the other hand, contributes to distinguishing more the effect of good predictions. In this scenario OptFTRL-C achieves an improvement of $16.8\%$ and $69.8\%$ over GPC for $\rho=0.1$ and $\rho=0.9$, respectively. In summary, OptFTRL-C leverages predictions without sacrificing its resilience to their inaccuracy, or to the costs’ adversity.

### v2

@@ V Numerical Example
Recall that OptFTRL-C was designed to take advantage of a prediction oracle that forecasts future cost functions with unknown accuracy. Theorem 1 then demonstrated that the average policy regret $\nicefrac{{\mathcal{R}_{T}}}{{T}}$ of OptFTRL-C always converges to 00, but does so faster for accurate predictions. We therefore plot the policy regret of OptFTRL-C when provided with either accurate or inaccurate predictions. The implementation code of the policies OptFTRL-C, GPC, and the benchmark $\pi^{\star}$, as well as the code necessary to reproduce all experiments, is available at the repository [47].
@@
In the accurate prediction case, we set $\tilde{c}_{t}(\cdot,\cdot)=c_{t}(\cdot,\cdot)$ with probability $\rho=0.9$. Otherwise, we set $\tilde{c}_{t}(\cdot,\cdot)$ to be uniformally random ($\boldsymbol{\alpha}_{t}\in[-1,1]^{2}$). Hence, $\rho$ represents the probability of correctly predicting $c_{t}(\cdot,\cdot)$, and we sample it at every slot. For inaccurate prediction, we set $\rho=0.1$. We compare with GPC, which was shown to outperform the classical $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ controllers in a wide range of situations.
@@
In scenario (b), we deploy an alternating cost function. Namely, $\boldsymbol{\alpha}_{t}$ alternates between $(1,1)$ and $(-0.5,-0.5)$ every $50$ steps. The disturbances are still $\boldsymbol{w}_{t}\!=\!(1,1),\forall t$. This alternating cost represents an adversarial fluctuation in the cost trajectory, and it is where the non-stochastic framework demonstrates its efficacy. Namely, this fluctuation is enough to violate the guarantees of $\mathcal{H}_{2}$ controllers, but at the same time, it is small in magnitude, rendering $\mathcal{H}_{\infty}$ controller overly pessimistic. It is worth noting that the fluctuation observed in Fig. $b$ is attributed to the update rule of GPC. This update directly modifies the decision variables $M_{t+1}$ based on the observed cost. In contrast, FTRL aggregates both past and future costs to determine $M_{t+1}$. With accurate predictions, this method does not induce much fluctuation as it foresees the upcoming small disturbance. Overall, OptFTRL-C with good prediction achieves an improvement of $32\%$ in $\nicefrac{{\mathcal{R}_{T}}}{{T}}$ value over GPC, while having a $13.4\%$ degradation when fed with the inaccurate oracle.
In scenario (c) we also deploy an alternating cost function but with different lower magnitudes. Namely, $\boldsymbol{\alpha}_{t}$ alternates between $(0.1,0.1)$ and $(-0.5,-0.5)$ every $50$ steps, and $w_{t}=(0.1,0.1)$. The goal of this scenario is to show that OptFTRL-C can have an advantage over GPC even regardless of the prediction quality through adapatablity to “easy” environments (i.e., environments with small gradients). In general, GPC performance takes a hit since its learning rate is tuned with the upper bound for the gradient (i.e., $\alpha=w=\sqrt{2},T=10^{3}$). The alternating frequency, on the other hand, contributes to distinguishing more the effect of good predictions. In this scenario OptFTRL-C achieves an improvement of $16.8\%$ and $69.8\%$ over GPC for $\rho=0.1$ and $\rho=0.9$, respectively. In summary, OptFTRL-C leverages predictions without sacrificing its resilience to their inaccuracy, or to the costs’ adversity.
@@ VI Conclusion
VII Acknowledgment
This work was supported by the Dutch National Growth Fund project “Future Network Services” and by the European Commission through Grant No. 101139270 “ORIGAMI”.